# Evaluating Rare Disease Diagnostic Performance in Symptom Checkers: A Synthetic Vignette Simulation Approach

### v2

@@ Abstract
Background: Symptom Checkers (SCs) provide medical information tailored to the user’s characteristics and symptoms. SC developers continuously improve their AI or algorithms, and to avoid unexpected performance degradation; they need to evaluate individual diseases’ diagnostic performance changes before deployment. However, sufficient evaluation data is difficult to obtain for rare diseases due to limited feedback from SC’s users. Creating dozens of clinical vignettes for each rare disease is also costly and impractical.
Objective: This study aims to propose a novel Synthetic Vignette Simulation Approach to evaluate diagnostic performance changes for individual rare diseases due to SC algorithm updates before deployment and to validate its effectiveness.
Methods: We used the disease-phenotype annotations from the Human Phenotype Ontology (HPO), a medical knowledge database for rare diseases. Using synthetic vignettes generated from HPO annotations, we simulated SC interviews and estimated the impact of algorithm updates on diagnostic performance in the real-world service. The proposed method’s effectiveness was evaluated retrospectively using R-squared by comparing the estimated values with the actual metrics changes reproduced in offline tests.
Results: The experiment included eight past SC algorithm updates for rare diseases. For algorithm updates targeting diseases with frequency information in HPO annotations (n=5), the R-squared coefficient for Recall@8 change was 0.831 ($p$-value: 0.031), and for Precision@8 change, it was 0.78 ($p$-value: 0.047). These results show that the proposed method can predict the performance change after deployment. Large prediction errors were shown for algorithm updates for diseases without frequency information in the HPO annotation (n=3), indicating that frequency information is essential for SC interview simulation. For all algorithm changes included in the experiment (n=8), the R-squared coefficient for Recall@8 change was 0.69 ($p$-value: 0.010), and for Precision@8 change, it was 0.16 ($p$-value: 0.318). We took approximately 2 hours per disease to map the Phenotypes in HPO Annotation to the SC symptoms.
Conclusions: Our proposed method enables evaluating the impact of SC algorithm changes on individual rare diseases before deployment. This evaluation is based on a publicly available medical knowledge database created by experts, ensuring transparency and explainability for stakeholders. Additionally, SC developers can efficiently improve diagnostic performance for rare diseases at a low cost. Our method could help SCs to enhance early diagnosis and treatment support for users.
@@ 1.2 Challenges in Evaluating Symptom Checker Performance for Rare Diseases: A Developer’s Perspective
SC developers continuously work on improving diagnostic performance, including for individual rare diseases. Each time they update the algorithm, they want to evaluate performance for specific diseases objectively. This evaluation ensures improvements in the target diseases and that diagnostic accuracy for other diseases does not decrease. Without sufficient performance evaluation, algorithm updates that degrade performance and go unnoticed can be released. For rare diseases, there is no established method to mitigate this risk. Generally, SC performance testing uses vignettes, but creating sufficient clinical vignettes for each disease is costly and impractical. Few specialists are capable of making these vignettes for rare diseases. Developers need more than 100 vignettes per disease to achieve a 95% confidence interval for sensitivity smaller than 0.1. The vignette set used in Hammoud et al.’s study [[Hammoud2024-hz]] covers 254 diseases with 400 vignettes in total, having a maximum of 5 vignettes per disease, which is insufficient for evaluating disease-specific performance. Some studies construct performance evaluation datasets based on case reports [[Miyachi2023-sr]], but case reports often cover rare instances and do not adequately represent typical cases. An ideal evaluation dataset would comprehensively cover typical cases across various age groups and genders, including early and late stages of disease onset. While SC developers can collect user feedback continuously to evaluate performance, the limited feedback for rare diseases makes data collection time-consuming. To address this issue, we propose a method for evaluating SC diagnostic performance for rare diseases using a medical knowledge database.
1.3 Simulating Real-World Symptom Checker Performance with Medical Knowledge Databases
A single medical professional cannot gain extensive clinical experience in rare diseases. This limitation has driven the development of medical knowledge databases to facilitate knowledge sharing and computational analysis. Orphanet [[Weinreich2008-mi]] and OMIM [[Hamosh2005-op]] are well-known, and a database of disease phenotype annotations coded using the Human Phenotype Ontology (HPO) [[Gargano2024-fc]], which systematizes abnormal clinical phenotypes in humans, is available. Disease phenotype annotations consist of clinical phenotypes that occur in diseases and their frequencies. We intended that synthetic vignettes generated from medical knowledge databases could serve as inputs for SCs. By simulating interviews with these vignettes, we aimed to estimate the impact of algorithm updates on service performance. Our proposed method has the following features:

### v3

@@ Abstract
Background: Symptom Checkers (SCs) provide medical information tailored to user symptoms. A critical challenge in SC development is preventing unexpected performance degradation for individual diseases, especially rare diseases, when updating algorithms. This risk stems from the lack of practical pre-deployment evaluation methods. For rare diseases, obtaining sufficient evaluation data from user feedback is difficult, and creating numerous clinical vignettes is costly and impractical.
Objective: To evaluate the impact of algorithm updates on the diagnostic performance for individual rare diseases before deployment, this study proposes and validates a novel Synthetic Vignette Simulation Approach. This approach aims to enable this essential evaluation efficiently and at a low cost.
Methods: To estimate the impact of algorithm updates, we generated synthetic vignettes from disease-phenotype annotations in the Human Phenotype Ontology (HPO), a publicly available knowledge base for rare diseases curated by experts. Using these vignettes, we simulated SC interviews to predict changes in diagnostic performance. The effectiveness of this approach was validated retrospectively by comparing the predicted changes with actual performance metrics (reproduced in offline tests) using the R-squared ($R^{2}$) coefficient.
Results: Our experiment, covering eight past algorithm updates for rare diseases, showed that the proposed method accurately predicted performance changes for diseases with phenotype frequency information in HPO (n=5). For these updates, we found a strong correlation for both Recall@8 change ($R^{2}=0.83$,$p=0.031$) and Precision@8 change ($R^{2}=0.78$,$p=0.047$). In contrast, updates for diseases without frequency data (n=3) resulted in large prediction errors, highlighting the critical role of this information. Mapping HPO phenotypes to SC symptoms required approximately 2 hours per disease.
Conclusions: Our proposed method enables the pre-deployment evaluation of SC algorithm changes for individual rare diseases. This evaluation is based on a publicly available medical knowledge database created by experts, ensuring transparency and explainability for stakeholders. Additionally, SC developers can efficiently improve diagnostic performance for rare diseases at a low cost.
@@ 1.2 Challenges in Evaluating Symptom Checker Performance for Rare Diseases: A Developer’s Perspective
SC developers continuously work on improving diagnostic performance, including for individual rare diseases. Each time they update the algorithm, they want to evaluate performance for specific diseases objectively. This evaluation ensures improvements in the target diseases and that diagnostic accuracy for other diseases does not decrease. Without sufficient performance evaluation, algorithm updates that degrade performance and go unnoticed can be released. For rare diseases, there is no established method to mitigate this risk. Generally, SC performance testing uses vignettes, but creating sufficient clinical vignettes for each disease is costly and impractical. Few specialists are capable of making these vignettes for rare diseases. Developers need more than 100 vignettes per disease to achieve a 95% confidence interval for sensitivity smaller than 0.1. The vignette set used in Hammoud et al.’s study [[Hammoud2024-hz]] covers 254 diseases with 400 vignettes in total, having a maximum of 5 vignettes per disease, which is insufficient for evaluating disease-specific performance. Some studies construct performance evaluation datasets based on case reports [[Miyachi2023-sr]], but case reports often cover rare instances and do not adequately represent typical cases. An ideal evaluation dataset would comprehensively cover typical cases across various age groups and genders, including early and late stages of disease onset. While SC developers can collect user feedback continuously to evaluate performance, the limited feedback for rare diseases makes data collection time-consuming.
1.3 Proposal: Simulating Real-World Symptom Checker Performance with Medical Knowledge Databases
To address this issue, we propose a method for evaluating SC diagnostic performance for rare diseases using a medical knowledge database. A single medical professional cannot gain extensive clinical experience in rare diseases. This limitation has driven the development of medical knowledge databases to facilitate knowledge sharing and computational analysis. Orphanet [[Weinreich2008-mi]] and OMIM [[Hamosh2005-op]] are well-known, and a database of disease phenotype annotations coded using the Human Phenotype Ontology (HPO) [[Gargano2024-fc]], which systematizes abnormal clinical phenotypes in humans, is available. Disease phenotype annotations consist of clinical phenotypes that occur in diseases and their frequencies. We intended that synthetic vignettes generated from medical knowledge databases could serve as inputs for SCs. By simulating interviews with these vignettes, we aimed to estimate the impact of algorithm updates on service performance. Our proposed method has the following features: